# Complex Wavelet Mutual Information Loss: A Multi-Scale Loss Function for Semantic Segmentation

### v1

@@ Abstract
Recent advancements in deep neural networks have significantly enhanced the performance of semantic segmentation. However, class imbalance and instance imbalance remain persistent challenges, where smaller instances and thin boundaries are often overshadowed by larger structures. To address the multiscale nature of segmented objects, various models have incorporated mechanisms such as spatial attention and feature pyramid networks. Despite these advancements, most loss functions are still primarily pixel-wise, while regional and boundary-focused loss functions often incur high computational costs or are restricted to small-scale regions. To address this limitation, we propose complex wavelet mutual information (CWMI) loss, a novel loss function that leverages mutual information from subband images decomposed by a complex steerable pyramid. The complex steerable pyramid captures features across multiple orientations and preserves structural similarity across scales. Meanwhile, mutual information is well-suited for capturing high-dimensional directional features and exhibits greater noise robustness. Extensive experiments on diverse segmentation datasets demonstrate that CWMI loss achieves significant improvements in both pixel-wise accuracy and topological metrics compared to state-of-the-art methods, while introducing minimal computational overhead. The code is available at https://anonymous.4open.science/r/CWMI-83B7/
@@ 2 Related Work
Semantic segmentation has seen tremendous advancements through deep learning architectures, with U-Net and its variants becoming a cornerstone of this field. The U-Net model, proposed by (Ronneberger et al., 2015), utilizes a symmetric encoder-decoder architecture with skip connections to preserve spatial details while capturing global context. Its success has inspired numerous adaptations in its convolutional blocks (Huang et al., 2019; Diakogiannis et al., 2020) and skip connections (Zhou et al., 2018; Chen et al., 2021). Inspired by the transformer model, the attention mechanism has also been incorporated into the U-Net structure, which has shown significant performance enhancement, as in Attention U-Net (Islam et al., 2020), TransUNet (Chen et al., 2021), etc. In this study, we compare our proposed CWMI loss using U-Net and Attention U-Net to test the generalization and superiority of CWMI, while the general idea is adaptable to other architectures as well.
@@ 2.1 Loss Functions for Semantic Segmentation
Several methods address instance imbalance dynamically during training but at the cost of computational efficiency. Topology-based approaches, such as persistent homology methods (Stucki et al., 2023; Oner et al., 2023), describe image topologies and identify critical pixels but are computationally expensive, with cubic complexity to image size. The clDice loss (Shit et al., 2021) employs a soft skeletonization algorithm to detect topological errors, primarily focusing on thin-boundary objects like retinal blood vessels. Similarly, Boundary Loss (Kervadec et al., 2019) and Hausdorff Distance Loss (Karimi & Salcudean, 2019) refine boundaries but incur significant computational overhead. Region Mutual Information (RMI) loss (Zhao et al., 2019) captures pixel interdependencies over regions, but struggles with scalability for large-scale regional analysis due to the high computation cost. These losses either prioritize small regions at the expense of global accuracy or require extensive computational resources, necessitating more efficient and balanced approaches.
@@ 3.2 Mutual Information in Complex Wavelet Domain
According to Equation 5, the ground truth $\mathbf{Y}$ and the prediction $\mathbf{P}$ are decomposed into subbands $\mathbf{Y}_{B_{n}}$ and $\mathbf{P}_{B_{n}}\in\mathbf{R}^{K\times H_{n-1}\times W_{n-1}}$ for each level $n\in[1,N]$. For each pixel $(x,y)$ at level $n$, the $K$-directional features are treated as $K$-dimensional random variables. The mutual information between $\mathbf{Y}_{B_{n}}$ and $\mathbf{P}_{B_{n}}$ is:
$I(\mathbf{Y}_{B_{n}};\mathbf{P}_{B_{n}})=H(\mathbf{Y}_{B_{n}})-H(\mathbf{Y}_{B_{n}}|\mathbf{P}_{B_{n}}),$ (7)
where $H(\mathbf{Y}_{B_{n}})$ is constant. Following (Zhao et al., 2019), we estimate the mutual information with:
@@
where $\mathbf{\Sigma}_{\mathbf{Y}_{B_{n}}}$ and $\mathbf{\Sigma}_{\mathbf{P}_{B_{n}}}$ are covariance matrices, and $\text{Cov}(\mathbf{Y}_{B_{n}},\mathbf{P}_{B_{n}})$ is the cross-covariance. For the complex field, this equation is extended with Hermitian transpose $H$:
@@
CE $0.751_{\pm 0.005}$ $0.850_{\pm 0.004}$ $2.23_{\pm 0.27}$ $0.477_{\pm 0.041}$ $1.21_{\pm 0.21}$ $0.754_{\pm 0.006}$ $0.852_{\pm 0.004}$ $1.69_{\pm 0.15}$ $0.558_{\pm 0.030}$ $1.23_{\pm 0.36}$
BCE $0.736_{\pm 0.007}$ $0.841_{\pm 0.005}$ $1.80_{\pm 0.21}$ $0.558_{\pm 0.029}$ $1.45_{\pm 0.03}$ $0.736_{\pm 0.014}$ $0.841_{\pm 0.010}$ $1.79_{\pm 0.03}$ $0.561_{\pm 0.011}$ $1.40_{\pm 0.12}$
Dice $0.767_{\pm 0.005}$ $0.861_{\pm 0.004}$ $1.38_{\pm 0.06}$ $0.605_{\pm 0.016}$ $0.91_{\pm 0.20}$ $0.767_{\pm 0.007}$ $0.861_{\pm 0.005}$ $1.53_{\pm 0.11}$ $0.580_{\pm 0.021}$ $0.81_{\pm 0.05}$
Focal $0.729_{\pm 0.006}$ $0.836_{\pm 0.005}$ $1.72_{\pm 0.06}$ $0.562_{\pm 0.014}$ $1.72_{\pm 0.23}$ $0.724_{\pm 0.004}$ $0.832_{\pm 0.003}$ $2.01_{\pm 0.19}$ $0.529_{\pm 0.009}$ $1.81_{\pm 0.14}$
Jaccard $0.766_{\pm 0.003}$ $0.861_{\pm 0.003}$ $1.36_{\pm 0.11}$ $0.615_{\pm 0.019}$ $0.89_{\pm 0.06}$ $0.764_{\pm 0.001}$ $0.860_{\pm 0.000}$ $1.36_{\pm 0.10}$ $0.609_{\pm 0.009}$ $0.90_{\pm 0.04}$
Tversky $0.764_{\pm 0.005}$ $0.860_{\pm 0.003}$ $1.30_{\pm 0.01}$ $0.622_{\pm 0.004}$ $1.07_{\pm 0.19}$ $0.765_{\pm 0.004}$ $0.860_{\pm 0.003}$ $1.28_{\pm 0.01}$ $0.623_{\pm 0.004}$ $1.01_{\pm 0.12}$
WCE $0.715_{\pm 0.006}$ $0.826_{\pm 0.003}$ $2.18_{\pm 0.53}$ $0.505_{\pm 0.055}$ $1.50_{\pm 0.23}$ $0.713_{\pm 0.006}$ $0.826_{\pm 0.005}$ $1.80_{\pm 0.04}$ $0.546_{\pm 0.014}$ $1.59_{\pm 0.15}$
ABW $0.637_{\pm 0.051}$ $0.769_{\pm 0.039}$ $2.56_{\pm 0.59}$ $0.415_{\pm 0.085}$ $2.83_{\pm 1.51}$ $0.639_{\pm 0.036}$ $0.769_{\pm 0.028}$ $2.62_{\pm 0.45}$ $0.410_{\pm 0.065}$ $2.88_{\pm 0.80}$
Skea-topo $0.658_{\pm 0.002}$ $0.785_{\pm 0.002}$ $2.34_{\pm 0.00}$ $0.449_{\pm 0.000}$ $1.62_{\pm 0.21}$ $0.657_{\pm 0.005}$ $0.784_{\pm 0.004}$ $2.42_{\pm 0.04}$ $0.446_{\pm 0.007}$ $1.47_{\pm 0.00}$
RMI $0.767_{\pm 0.009}$ $0.861_{\pm 0.007}$ $1.46_{\pm 0.17}$ $0.587_{\pm 0.038}$ $\underline{0.82_{\pm 0.06}}$ $0.763_{\pm 0.009}$ $0.858_{\pm 0.007}$ $1.46_{\pm 0.17}$ $0.577_{\pm 0.031}$ $0.94_{\pm 0.09}$
clDice $0.706_{\pm 0.023}$ $0.819_{\pm 0.018}$ $2.26_{\pm 0.41}$ $0.476_{\pm 0.068}$ $1.56_{\pm 0.28}$ $0.690_{\pm 0.014}$ $0.807_{\pm 0.010}$ $1.97_{\pm 0.15}$ $0.498_{\pm 0.023}$ $1.82_{\pm 0.15}$
CWMI-Real $\underline{0.774_{\pm 0.005}}$ $\underline{0.866_{\pm 0.004}}$ $\underline{1.19_{\pm 0.10}}$ $\underline{0.634_{\pm 0.018}}$ $0.91_{\pm 0.20}$ $\underline{0.775_{\pm 0.005}}$ $\underline{0.867_{\pm 0.004}}$ $\mathbf{1.16_{\pm 0.07}}$ $\mathbf{0.638_{\pm 0.010}}$ $\underline{0.80_{\pm 0.09}}$
CWMI $\mathbf{0.779_{\pm 0.004}}$ $\mathbf{0.869_{\pm 0.003}}$ $\mathbf{1.16_{\pm 0.09}}$ $\mathbf{0.640_{\pm 0.005}}$ $\mathbf{0.69_{\pm 0.04}}$ $\mathbf{0.778_{\pm 0.009}}$ $\mathbf{0.869_{\pm 0.007}}$ $\underline{1.19_{\pm 0.15}}$ $\underline{0.637_{\pm 0.034}}$ $\mathbf{0.78_{\pm 0.08}}$
@@
CE $0.808_{\pm 0.009}$ $0.889_{\pm 0.007}$ $0.99_{\pm 0.02}$ $0.686_{\pm 0.026}$ $3.13_{\pm 0.65}$ $0.811_{\pm 0.009}$ $0.889_{\pm 0.005}$ $1.01_{\pm 0.10}$ $0.690_{\pm 0.028}$ $3.44_{\pm 0.03}$
BCE $0.682_{\pm 0.038}$ $0.798_{\pm 0.031}$ $1.82_{\pm 0.57}$ $0.309_{\pm 0.154}$ $10.87_{\pm 5.26}$ $0.699_{\pm 0.045}$ $0.814_{\pm 0.032}$ $1.48_{\pm 0.55}$ $0.398_{\pm 0.137}$ $9.77_{\pm 5.23}$
Dice $0.816_{\pm 0.019}$ $0.894_{\pm 0.014}$ $0.94_{\pm 0.06}$ $0.699_{\pm 0.019}$ $3.01_{\pm 0.87}$ $0.817_{\pm 0.019}$ $0.895_{\pm 0.013}$ $0.91_{\pm 0.15}$ $0.704_{\pm 0.031}$ $2.70_{\pm 0.59}$
Focal $0.583_{\pm 0.080}$ $0.715_{\pm 0.075}$ $2.31_{\pm 0.20}$ $0.137_{\pm 0.105}$ $16.50_{\pm 9.43}$ $0.588_{\pm 0.032}$ $0.719_{\pm 0.028}$ $2.15_{\pm 0.25}$ $0.151_{\pm 0.037}$ $23.18_{\pm 5.86}$
Jaccard $0.803_{\pm 0.024}$ $0.886_{\pm 0.016}$ $1.02_{\pm 0.12}$ $0.675_{\pm 0.050}$ $3.30_{\pm 0.63}$ $0.812_{\pm 0.010}$ $0.892_{\pm 0.006}$ $0.92_{\pm 0.03}$ $0.693_{\pm 0.019}$ $3.10_{\pm 0.54}$
Tversky $0.804_{\pm 0.024}$ $0.887_{\pm 0.016}$ $0.97_{\pm 0.13}$ $0.670_{\pm 0.039}$ $3.18_{\pm 0.70}$ $0.814_{\pm 0.024}$ $0.892_{\pm 0.016}$ $0.93_{\pm 0.07}$ $0.698_{\pm 0.047}$ $3.34_{\pm 1.26}$
WCE $0.829_{\pm 0.010}$ $0.902_{\pm 0.007}$ $0.83_{\pm 0.07}$ $0.721_{\pm 0.008}$ $\mathbf{2.26_{\pm 0.21}}$ $0.823_{\pm 0.020}$ $0.898_{\pm 0.014}$ $0.87_{\pm 0.13}$ $0.705_{\pm 0.026}$ $\mathbf{2.47_{\pm 0.61}}$
ABW $0.765_{\pm 0.011}$ $0.860_{\pm 0.008}$ $1.43_{\pm 0.03}$ $0.621_{\pm 0.033}$ $3.65_{\pm 0.72}$ $0.766_{\pm 0.011}$ $0.862_{\pm 0.008}$ $1.49_{\pm 0.11}$ $0.614_{\pm 0.018}$ $3.20_{\pm 0.58}$
Skea-topo $0.788_{\pm 0.008}$ $0.876_{\pm 0.005}$ $1.26_{\pm 0.04}$ $0.658_{\pm 0.002}$ $2.99_{\pm 0.76}$ $0.784_{\pm 0.016}$ $0.873_{\pm 0.010}$ $1.30_{\pm 0.09}$ $0.645_{\pm 0.022}$ $3.36_{\pm 0.61}$
RMI $\underline{0.839_{\pm 0.003}}$ $\underline{0.908_{\pm 0.002}}$ $0.82_{\pm 0.06}$ $\underline{0.733_{\pm 0.003}}$ $2.50_{\pm 0.13}$ $\underline{0.842_{\pm 0.001}}$ $\underline{0.910_{\pm 0.001}}$ $0.83_{\pm 0.04}$ $\underline{0.732_{\pm 0.006}}$ $2.58_{\pm 0.61}$
clDice $0.818_{\pm 0.010}$ $0.895_{\pm 0.007}$ $0.89_{\pm 0.05}$ $0.696_{\pm 0.020}$ $2.70_{\pm 0.17}$ $0.803_{\pm 0.015}$ $0.886_{\pm 0.011}$ $0.96_{\pm 0.07}$ $0.687_{\pm 0.025}$ $3.10_{\pm 0.42}$
CWMI-Real $0.835_{\pm 0.008}$ $0.905_{\pm 0.004}$ $\underline{0.80_{\pm 0.03}}$ $0.724_{\pm 0.010}$ $2.74_{\pm 0.60}$ $0.842_{\pm 0.022}$ $0.909_{\pm 0.015}$ $\underline{0.79_{\pm 0.06}}$ $0.724_{\pm 0.038}$ $2.76_{\pm 0.76}$
CWMI $\mathbf{0.845_{\pm 0.013}}$ $\mathbf{0.911_{\pm 0.010}}$ $\mathbf{0.73_{\pm 0.04}}$ $\mathbf{0.745_{\pm 0.017}}$ $\underline{2.47_{\pm 0.42}}$ $\mathbf{0.844_{\pm 0.007}}$ $\mathbf{0.911_{\pm 0.005}}$ $\mathbf{0.75_{\pm 0.11}}$ $\mathbf{0.737_{\pm 0.015}}$ $\underline{2.57_{\pm 0.55}}$

### v2

@@ Abstract
Recent advancements in deep neural networks have significantly enhanced the performance of semantic segmentation. However, class imbalance and instance imbalance remain persistent challenges, where smaller instances and thin boundaries are often overshadowed by larger structures. To address the multiscale nature of segmented objects, various models have incorporated mechanisms such as spatial attention and feature pyramid networks. Despite these advancements, most loss functions are still primarily pixel-wise, while regional and boundary-focused loss functions often incur high computational costs or are restricted to small-scale regions. To address this limitation, we propose the complex wavelet mutual information (CWMI) loss, a novel loss function that leverages mutual information from subband images decomposed by a complex steerable pyramid. The complex steerable pyramid captures features across multiple orientations and preserves structural similarity across scales. Meanwhile, mutual information is well-suited to capturing high-dimensional directional features and offers greater noise robustness. Extensive experiments on diverse segmentation datasets demonstrate that CWMI loss achieves significant improvements in both pixel-wise accuracy and topological metrics compared to state-of-the-art methods, while introducing minimal computational overhead. Our code is available at https://github.com/lurenhaothu/CWMI
@@ 2 Related Work
Semantic segmentation has seen tremendous advancements through deep learning architectures, with U-Net and its variants becoming a cornerstone of this field. The U-Net model, proposed by (Ronneberger et al., 2015), utilizes a symmetric encoder-decoder architecture with skip connections to preserve spatial details while capturing global context. Its success has inspired numerous adaptations in its convolutional blocks (Huang et al., 2019; Diakogiannis et al., 2020) and skip connections (Zhou et al., 2018; Chen et al., 2021). Inspired by the transformer model, the attention mechanism has also been incorporated into the U-Net structure, which has shown significant performance enhancement, as in Attention U-Net (Islam et al., 2020), TransUNet (Chen et al., 2021), Vision Mamba UNet (Ruan et al., 2024), etc. In this study, we compare our proposed CWMI loss using U-Net and Attention U-Net to test the generalization and superiority of CWMI, while the general idea is adaptable to other architectures as well.
@@ 2.1 Loss Functions for Semantic Segmentation
Several methods address instance imbalance dynamically during training, but at the cost of computational efficiency. Topology-based approaches, such as persistent homology methods (Stucki et al., 2023; Oner et al., 2023), describe image topologies and identify critical pixels but are computationally expensive, with cubic complexity to image size. The clDice loss (Shit et al., 2021) employs a soft skeletonization algorithm to detect topological errors, primarily focusing on thin-boundary objects like retinal blood vessels. Similarly, Boundary Loss (Kervadec et al., 2019) and Hausdorff Distance Loss (Karimi & Salcudean, 2019) refine boundaries but incur significant computational overhead. Region Mutual Information (RMI) loss (Zhao et al., 2019) captures pixel interdependencies over regions, but struggles with scalability for large-scale regional analysis due to the high computation cost. These losses either prioritize small regions at the expense of global accuracy or require extensive computational resources, necessitating more efficient and balanced approaches.
@@ 3.2 Mutual Information in Complex Wavelet Domain
According to Equation 5, the ground truth $\mathbf{Y}$ and the prediction $\mathbf{P}$ are decomposed into subbands $\mathbf{Y}_{B_{n}}$ and $\mathbf{P}_{B_{n}}\in\mathbf{R}^{K\times H_{n-1}\times W_{n-1}}$ for each level $n\in[1,N]$. For each pixel $(x,y)$ at level $n$, the $K$-directional features are treated as $K$-dimensional random variables.
For mutual information approximation, several studies, such as MINE (Belghazi et al., 2018) and Deep InfoMax (Hjelm et al., 2018), employ neural network-based estimators to produce tight lower bounds. While these methods are theoretically well-founded, they introduce additional training overhead, which increases algorithmic complexity and may hinder the efficiency of the loss function. Therefore, in this work, we adopt the mutual information estimation approach proposed by (Zhao et al., 2019):
@@
where $\mathbf{\Sigma}_{\mathbf{Y}_{B_{n}}}$ and $\mathbf{\Sigma}_{\mathbf{P}_{B_{n}}}$ are covariance matrices, and $\text{Cov}(\mathbf{Y}_{B_{n}},\mathbf{P}_{B_{n}})$ is the cross-covariance.
When the ground truth and predictions are decomposed into complex subbands, this equation is extended with Hermitian transpose $H$:
@@
where the covariance and cross-covariance are calculated as:
$\mathbf{\Sigma}_{\mathbf{Y}_{B_{n}}}=E[(\mathbf{Y}_{B_{n}}-E[\mathbf{Y}_{B_{n}}])(\mathbf{Y}_{B_{n}}-E[\mathbf{Y}_{B_{n}}])^{H}]$ (10)
$\mathbf{\Sigma}_{\mathbf{P}_{B_{n}}}=E[(\mathbf{P}_{B_{n}}-E[\mathbf{P}_{B_{n}}])(\mathbf{P}_{B_{n}}-E[\mathbf{P}_{B_{n}}])^{H}]$ (11)
$\text{Cov}(\mathbf{Y}_{B_{n}},\mathbf{P}_{B_{n}})=E[(\mathbf{Y}_{B_{n}}-E[\mathbf{Y}_{B_{n}}])(\mathbf{P}_{B_{n}}-E[\mathbf{P}_{B_{n}}])^{H}]$ (12)
@@
CE .$753_{\pm.005}.$$851_{\pm.004}.$ $2.082_{\pm.270}$$500_{\pm.039}.$ $1.143_{\pm.176}$  .$751_{\pm.009}.$$850_{\pm.007}.$ $1.914_{\pm.361}$$525_{\pm.056}.$ $1.178_{\pm.261}$
BCE .$736_{\pm.006}.$$841_{\pm.004}.$ $1.756_{\pm.156}$$561_{\pm.020}.$ $1.413_{\pm.088}$  .$735_{\pm.010}.$$840_{\pm.007}.$ $1.835_{\pm.150}$$553_{\pm.023}.$ $1.417_{\pm.224}$
Dice .$767_{\pm.004}.$$862_{\pm.003}.$ $1.406_{\pm.085}$$604_{\pm.017}.$$919_{\pm.154}.$  .$768_{\pm.005}.$$862_{\pm.004}.$ $1.507_{\pm.086}$$583_{\pm.016}.$$\mathbf{.804_{\pm.072}}$
Focal .$728_{\pm.007}.$$835_{\pm.005}.$ $1.751_{\pm.078}$$556_{\pm.016}.$ $1.725_{\pm.225}$  .$725_{\pm.005}.$$833_{\pm.003}.$ $1.911_{\pm.189}$$540_{\pm.019}.$ $1.803_{\pm.107}$
Jaccard .$766_{\pm.004}.$$861_{\pm.003}.$ $1.357_{\pm.114}$$614_{\pm.019}.$$904_{\pm.087}.$  .$762_{\pm.003}.$$858_{\pm.002}.$ $1.360_{\pm.079}$$609_{\pm.007}.$ $1.000_{\pm.125}$
Tversky .$765_{\pm.004}.$$861_{\pm.003}.$ $1.328_{\pm.077}$$617_{\pm.017}.$$990_{\pm.183}.$  .$765_{\pm.003}.$$860_{\pm.002}.$ $1.303_{\pm.057}$$621_{\pm.009}.$$969_{\pm.101}.$
WCE .$714_{\pm.005}.$$825_{\pm.003}.$ $1.987_{\pm.398}$$525_{\pm.042}.$ $1.575_{\pm.250}$  .$713_{\pm.005}.$$825_{\pm.004}.$ $1.818_{\pm.034}$$543_{\pm.009}.$ $1.613_{\pm.192}$
ABW .$605_{\pm.066}.$$739_{\pm.057}.$ $3.637_{\pm 1.703}$$299_{\pm.187}.$ $4.605_{\pm 2.926}$  .$616_{\pm.072}.$$749_{\pm.063}.$ $3.077_{\pm 1.303}$$366_{\pm.142}.$ $4.045_{\pm 3.220}$
Skea-topo .$572_{\pm.149}.$$671_{\pm.197}.$ $3.822_{\pm 2.560}$$300_{\pm.259}.$ $1.618_{\pm.207}$  .$602_{\pm.095}.$$729_{\pm.096}.$ $3.759_{\pm 2.324}$$306_{\pm.243}.$ $4.889_{\pm 5.928}$
RMI .$764_{\pm.006}.$$859_{\pm.005}.$ $1.443_{\pm.207}$$587_{\pm.037}.$$943_{\pm.130}.$  .$764_{\pm.008}.$$859_{\pm.006}.$ $1.404_{\pm.124}$$585_{\pm.028}.$$958_{\pm.170}.$
clDice .$706_{\pm.014}.$$819_{\pm.011}.$ $2.049_{\pm.351}$$502_{\pm.052}.$ $1.489_{\pm.218}$  .$701_{\pm.022}.$$815_{\pm.016}.$ $1.887_{\pm.161}$$518_{\pm.034}.$ $1.720_{\pm.244}$
Sensitive .$763_{\pm.006}.$$858_{\pm.005}.$ $1.543_{\pm.069}$$576_{\pm.023}.$$948_{\pm.143}.$  .$757_{\pm.006}.$$854_{\pm.004}.$ $1.822_{\pm.313}$$527_{\pm.049}.$ $1.016_{\pm.139}$
CWMI-Real .$\underline{.776_{\pm.004}}^{**}$$\underline{.867_{\pm.003}}^{**}$ $\underline{1.205_{\pm.059}}^{**}$$\underline{.634_{\pm.013}}^{*}$$\underline{.807_{\pm.117}}^{*}$  .$\underline{.775_{\pm.004}}^{**}$$\underline{.867_{\pm.003}}^{**}$ $\underline{1.193_{\pm.070}}^{**}$$\underline{.634_{\pm.014}}^{*}$$824_{\pm.103}.$
CWMI .$\mathbf{.778_{\pm.004}}^{***}$$\mathbf{.869_{\pm.003}}^{***}$ $\mathbf{1.162_{\pm.068}}^{***}$$\mathbf{.638_{\pm.015}}^{*}$$\mathbf{.739_{\pm.095}}^{**}$  .$\mathbf{.777_{\pm.006}}^{**}$$\mathbf{.868_{\pm.005}}^{**}$ $\mathbf{1.162_{\pm.085}}^{**}$$\mathbf{.639_{\pm.019}}^{*}$$\underline{.807_{\pm.069}}$
@@
CE .$637_{\pm.204}.$$724_{\pm.192}.$$933_{\pm.043}.$$400_{\pm.333}.$ $6.633_{\pm 4.654}$  .$640_{\pm.209}.$$725_{\pm.195}.$$906_{\pm.107}.$$406_{\pm.327}.$ $7.914_{\pm 6.637}$
BCE .$614_{\pm.226}.$$702_{\pm.213}.$ $1.032_{\pm.152}$$355_{\pm.355}.$ $4.165_{\pm 1.929}$  .$619_{\pm.222}.$$708_{\pm.208}.$$986_{\pm.139}.$$365_{\pm.358}.$ $4.354_{\pm 1.869}$
Dice .$632_{\pm.201}.$$719_{\pm.189}.$$887_{\pm.080}.$$396_{\pm.322}.$ $7.004_{\pm 4.670}$  .$636_{\pm.197}.$$723_{\pm.187}.$$954_{\pm.056}.$$396_{\pm.322}.$ $7.335_{\pm 4.843}$
Focal .$581_{\pm.253}.$$672_{\pm.241}.$ $1.093_{\pm.103}$$330_{\pm.356}.$ $4.447_{\pm 2.023}$  .$581_{\pm.252}.$$672_{\pm.240}.$ $1.043_{\pm.109}$$336_{\pm.369}.$ $5.040_{\pm 2.802}$
Jaccard .$637_{\pm.205}.$$723_{\pm.193}.$$947_{\pm.076}.$$397_{\pm.335}.$ $6.979_{\pm 4.511}$  .$638_{\pm.202}.$$725_{\pm.190}.$$888_{\pm.076}.$$406_{\pm.331}.$ $6.758_{\pm 4.431}$
Tversky .$634_{\pm.205}.$$722_{\pm.193}.$$953_{\pm.031}.$$394_{\pm.332}.$ $6.705_{\pm 4.398}$  .$627_{\pm.195}.$$717_{\pm.186}.$$924_{\pm.102}.$$384_{\pm.314}.$ $7.521_{\pm 4.493}$
WCE .$829_{\pm.010}.$$902_{\pm.007}.$$832_{\pm.075}.$$721_{\pm.008}.$ $\underline{2.264_{\pm.210}}$  .$823_{\pm.016}.$$898_{\pm.011}.$$873_{\pm.097}.$$705_{\pm.023}.$ $\underline{2.590_{\pm.602}}$
ABW .$760_{\pm.012}.$$857_{\pm.008}.$ $1.468_{\pm.079}$$619_{\pm.027}.$ $3.527_{\pm.508}$  .$762_{\pm.019}.$$858_{\pm.013}.$ $1.472_{\pm.103}$$614_{\pm.037}.$ $3.623_{\pm.887}$
Skea-topo .$788_{\pm.008}.$$876_{\pm.005}.$ $1.259_{\pm.042}$$658_{\pm.002}.$ $2.985_{\pm.765}$  .$784_{\pm.016}.$$873_{\pm.010}.$ $1.299_{\pm.086}$$645_{\pm.022}.$ $3.363_{\pm.609}$
RMI .$\underline{.839_{\pm.011}}$$\underline{.907_{\pm.007}}$$820_{\pm.082}.$$726_{\pm.024}.$ $2.644_{\pm.372}$  .$835_{\pm.015}.$$905_{\pm.010}.$$822_{\pm.083}.$$720_{\pm.039}.$ $2.701_{\pm.366}$
clDice .$816_{\pm.017}.$$894_{\pm.012}.$$920_{\pm.107}.$$698_{\pm.026}.$ $2.809_{\pm.476}$  .$800_{\pm.013}.$$884_{\pm.009}.$ $1.004_{\pm.113}$$684_{\pm.026}.$ $3.146_{\pm.286}$
Sensitive .$821_{\pm.014}.$$897_{\pm.009}.$$941_{\pm.101}.$$694_{\pm.033}.$ $2.916_{\pm.394}$  .$824_{\pm.018}.$$899_{\pm.011}.$$880_{\pm.108}.$$702_{\pm.041}.$ $2.757_{\pm.398}$
CWMI-Real .$838_{\pm.014}.$$907_{\pm.010}.$$\underline{.798_{\pm.068}}$$\underline{.727_{\pm.027}}$ $2.813_{\pm.653}$  .$\underline{.842_{\pm.022}}$$\underline{.909_{\pm.015}}$$\underline{.788_{\pm.063}}$$\underline{.724_{\pm.038}}$ $2.758_{\pm.764}$
CWMI .$\mathbf{.843_{\pm.016}}$$\mathbf{.910_{\pm.011}}$$\mathbf{.761_{\pm.081}}^{*}$$\mathbf{.735_{\pm.026}}$ $\mathbf{2.569_{\pm.466}}$  .$\mathbf{.844_{\pm.007}}$$\mathbf{.911_{\pm.005}}$$\mathbf{.755_{\pm.106}}$$\mathbf{.737_{\pm.015}}$ $\mathbf{2.569_{\pm.547}}$